The points D, E and F divide the sides BC, CA and AB of a triangle in the ratio 1:4, 3:2 and 3:7 respectively. Show that there exists a point G such that the sum of the vectors AD, BE, CF is parallel to CG. What is the ratio in which G divides AB?


1
Expert's answer
2021-01-28T19:04:12-0500

In triangle ABC, let A is zero vector.

B is "\overrightarrow{b}" and C is "\overrightarrow{c}" . D divides BC in 1:4 thus D is "\frac{\overrightarrow{c} + 4\overrightarrow{b}}{5}" . Similarly E is "\frac{2\overrightarrow{c}}{5}." F is "\frac{3\overrightarrow{b}}{10}" .


Now adding AD +BE+CF, we get


"\frac{\overrightarrow{c}+\overrightarrow{4b}}{5} - \overrightarrow{0} + \frac{\overrightarrow{2c}}{5} - \overrightarrow{b} + \frac{\overrightarrow{3b}}{10} - \overrightarrow{c}"


"=\frac{\overrightarrow{b}}{10} - \frac{\overrightarrow{2c}}{5}"



Now we know that point G divides the AB so that let it's divides in ratio 1:k then "\overrightarrow{G} = \frac{\overrightarrow{b}}{k+1}" and


"\overrightarrow{CG} = \frac{\overrightarrow{b}}{k+1} - \overrightarrow{c}" .


Now for k=3, "\overrightarrow{CG} = \frac{\overrightarrow{b}}{4} - \overrightarrow{c} ."


(2/5)"(\overrightarrow{CG}) = \frac{\overrightarrow{b}}{10} - \frac{\overrightarrow{2c}}{5}" .


We can see that addition of AD +BE +CF is equals to "\frac{2\overrightarrow{CG}}{5}." Hence it proves that the above sum is parallel to CG and also G divides the AB in 1:3 .